1. Given a 2-for-15 rights issue, a current market price of $13.50 and subscription price of $12, this implies a theoretical rights price of:

2. Given a current market price of $7.20, a subscription price of $5.50 and a theoretical rights price of $1.36, this implies a:

3. Given a 2-for-17 rights issue, a current market price of $7 and a subscription price of $6.50, the implied ex-rights price is:
Answer #1

Q. 1). Solution :-

Theoretical Ex- rights price = (Existing shares * market price per share + Number of right issue shares * Subscription price) / (Existing shares + Number of right issue shares)

= (15 * 13.50 + 2 * 12) / (15 + 2)

= (202.50 + 24.00) / 17

= 226.50 / 17

= $ 13.32

Theoretical rights price = Theoretical Ex- rights price - Subscription price of right issue.

= 13.32 - 12

= $ 1.32

Conclusion :- Theoretical right price = $ 1.32

Q. 3). Solution :-

Theoretical Ex- right price = (Existing shares * market price per share + Number of right issue shares * Subscription price) / (Existing shares + Number of right issue shares)

= (17 * 7 + 2 * 6.50) / (17 + 2)

= (119 + 13) / 19

= 132 / 19

= 6.95 (approx).

Conclusion :- Theoretical Ex-right price = $ 6.95 (approx).